- Some work for Y arm alignment.
Transition from morning work (*:responsible person)
Report of work (*:responsible person)
- MC was aligned and chamber doors was closed. MC is ready for IFO alignnment.
- CAL: Noise hunting. Cutting 5pin of D-SUB cable made 30dB improvement.
- PEM: Ondotori setting was finished. We will be added 10 more.
- AOS: Done finalization of Y front. Will work on ETMX this week. Needs heavy work inside the ETMX chamber.
Work plan (*:responsible person)
- Type B: We will do hight adjustments and some tuning for BS or other VIS.
- Type A: Needs work at ITMY today.
- DGS: We will separate the current rack at ETMX to the second one. It takes whole of this week.
- DGS: RFM ID will be changed at ETMX and ETMY.